# Low-Altitude ISAC with Rotatable Active and Passive Arrays

### v1

@@ IV Simulation Results
Figure 2: Sum rate versus maximum transmit power under $p^{\text{R}}=p^{\text{B}}=0$ with $N=6\times 6$, $M=2\times 2$ and $K=2$.
@@
Figure 3: Sum rate versus maximum transmit power under $p^{\text{R}}=p^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 2$ and $K=2$.
Figs. 2 and 3 illustrate the sum rate versus the BS transmit power under the sensing beampattern MSE constraint for a $2\times 2$ transmit array, two communication users, and two sensing directions. In the half-space isotropic case (Fig. 2), all schemes benefit from increasing power, showing roughly linear growth at low–to–moderate SNR and gradual saturation when multiuser interference and the sensing constraint become dominant. Across the entire power range, the proposed joint BS–RIS rotation (“Rot-BS+Rot-RIS”) consistently achieves the highest sum rate, followed by the partially rotatable schemes. At the same time, the non-RIS baselines perform worse due to their reduced effective aperture. However, the RIS array rotations seem not to be effective in the half-space isotropic case, because rotations in this case cannot reap the directivity gains; one can say that isotropic elements avoid severe signal degradation in side-lobe directions. With a directional half-space pattern (Fig. 3, element directivity factor $2$), all curves shift upward, confirming the SNR and throughput gains brought by element directivity. At the same time, the rate gap between Rot-BS+Rot-RIS and the benchmark schemes becomes more pronounced, as the proposed design better aligns the main lobes with user directions and steers reduced-gain regions toward the sensing directions, thereby alleviating the performance penalty of the sensing constraint. This indicates that geometric adaptation, especially when both active and passive arrays are rotatable, becomes increasingly valuable in the interference-limited regime and practical BS antenna or RIS element radiation patterns.
@@
Figure 4: Sum rate versus the number of BS antennas under $p^{\text{R}}=p^{\text{B}}=0$ with $N=6\times 6$, $M_{x}=1$ $P=30$ dBm, and $K=3$.
@@
Figure 5: Sum rate versus the number of BS antennas under $p^{\text{R}}=p^{\text{B}}=2$ with $N=6\times 6$, $M_{x}=1$ $P=30$ dBm, and $K=3$.
Figs. 4 and 5 examine the impact of the number of transmit antennas for a $1\times M_{y}$ BS array with three users. In the isotropic case (Fig. 4), the sum rate of all schemes increases with $M_{y}$, since a larger array aperture provides stronger array gain and finer spatial resolution to separate the three users while shaping the sensing beampattern. Schemes with only BS rotation or only RIS rotation still benefit from adding antennas, but their curves remain noticeably below the fully rotatable configuration, indicating that partial geometric adaptation cannot fully unlock the potential of a large array in the presence of both multiuser interference and sensing requirements. The non-RIS baselines lie at the bottom in all $M_{y}$, highlighting the importance of the extra reflecting aperture with more spatial resources. In the directional case (Fig. 5), all curves are further elevated and the advantage of Rot-BS+Rot-RIS becomes more marked. The directivity amplifies the relative advantage of the joint BS–RIS rotation: by aligning the main lobe of the transmit array and the dominant reflecting directions of the RIS with the user angles while steering reduced-gain regions toward the sensing directions, the proposed design can more efficiently decouple communication beams from sensing beams, leading to a larger rate gain over the partially rotatable and fully fixed benchmarks. In summary, combining array scaling with geometric adaptation is effective for one-dimensional arrays in ISAC systems, especially with directional arrays with practical radiation patterns.
@@
Figure 6: Sum rate versus the number of communication users under $p^{\text{R}}=p^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 4$, and $P=22$ dBm.
@@
Figure 7: Sum rate versus the number of communication users under $p^{\text{R}}=p^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 2$, and $P=30$ dBm.
Figs. 6 investigate user scaling for a $2\times 4$ BS array with fixed transmit power $P_{\mathrm{B}}=22$ dBm. The sum rate increases with the number of users due to multiuser multiplexing gains, but the growth gradually slows as the system becomes more interference-limited and the available spatial degrees of freedom must be shared between user separation and sensing beampattern shaping. The Rot-BS+Rot-RIS scheme consistently delivers the highest sum rate, demonstrating that joint geometry control is especially beneficial in the multiuser regime, whereas partial rotation or fixed architectures suffer from stronger interference and tighter sensing constraints.
Fig. 7 illustrates the impact of the maximum BS rotation range $\zeta$ on the achievable sum rate for different RIS array rotation ranges. As $\zeta$ increases, the sum rate of all schemes improves and gradually saturates, since a larger rotation range provides more geometric degrees of freedom to steer the BS beams toward the users and away from interference directions. Across the whole range of $\zeta$, both rotatable-RIS configurations significantly outperform the fixed-IRS and no-IRS baselines, confirming that jointly adapting the BS orientation and the IRS reflection direction is highly beneficial. Moreover, the curves for “rot-IRS $90^{\circ}$” and “rot-IRS $45^{\circ}$” almost coincide once $\zeta$ exceeds $30^{\circ}$, indicating that a moderate IRS rotation capability already captures most of the gain brought by passive array rotation, while the BS rotation range remains the dominant factor in leveraging geometric flexibility.
@@ V Conclusion
This paper studied a low-altitude ISAC system with cooperative rotatable active and passive arrays. We considered a downlink transmission where a BS equipped with an active RA serves multiple ground users and senses low-altitude targets in monostatic mode, assisted by a rotatable RIS, i.e., passive RA. To capture the impact of 3D array rotations, we developed a rotation-aware geometry-based multipath channel model in which BS–user, BS–RIS, and RIS–user links are parameterized by path-dependent 3D angles and complex gains, and array rotations jointly affect steering vectors and direction-dependent element gains. On this basis, we formulated a unified design that maximizes the multiuser sum rate under a transmit-power budget, RIS unit-modulus constraints, mechanical rotation limits, and a sensing beampattern MSE constraint. To tackle the resulting highly coupled and non-convex problem, we proposed a penalty-based AO framework that alternately updates the BS ISAC precoder, RIS phase-shift vector, and BS/RIS rotation angles. The BS precoding block is solved via a QT– and MM-aided method, leading to a tractable QCQP; the RIS block is handled on the complex-circle manifold using an RGD algorithm; and the rotation block is optimized via PGD with Barzilai–Borwein step sizes and Armijo backtracking. Extensive simulations under LAWN-inspired 3D geometries showed that the proposed jointly rotatable BS–RIS architecture achieves significant sum-rate gains over fixed or partially rotatable baselines while satisfying stringent sensing MSE requirements, with the advantages further amplified with directional antennas.

### v2

@@ IV-A Simulation Setup and Benchmarks
Figure 2: Weighted ISAC utility versus maximum transmit power under $b^{\text{R}}=b^{\text{B}}=0$ with $\rho=10$, $N=6\times 6$, $M=2\times 2$ and $K=2$.
@@
Figure 3: Weighted ISAC utility versus maximum transmit power under $b^{\text{R}}=b^{\text{B}}=2$ with $\rho=10$, $N=6\times 6$, $M=2\times 2$ and $K=2$.
@@
Figure 4: Sum rate versus maximum transmit power under $b^{\text{R}}=b^{\text{B}}=0$ with $\rho=10$, $N=6\times 6$, $M=2\times 2$ and $K=2$.
@@
Figure 5: Sum rate versus maximum transmit power under $b^{\text{R}}=b^{\text{B}}=2$ with $\rho=10$, $N=6\times 6$, $M=2\times 2$ and $K=2$.
@@
Figure 6: Beampattern NMSE versus maximum transmit power under $b^{\text{R}}=b^{\text{B}}=0$ with $\rho=10$, $N=6\times 6$, $M=2\times 2$ and $K=2$.
@@
Figure 7: Beampattern NMSE versus maximum transmit power under $b^{\text{R}}=b^{\text{B}}=2$ with $\rho=10$, $N=6\times 6$, $M=2\times 2$ and $K=2$.
IV-B Impact of Transmit Power
Figs. 2-7 evaluate the impact of the BS transmit-power budget $P_{\text{B}}$ under $M=2\times 2$, $N=6\times 6$, and $K=2$. Figs. 2-3 plot the weighted utility $\mathcal{U}$, while Figs. 4-5 and Figs. 6-7 plot the associated sum rate and NMSE, respectively.
Half-space isotropic elements ($b=0$): From Figs. 2, 4, and 6, increasing $P_{\text{B}}$ monotonically improves the achievable utility and sum rate for all schemes, as higher power boosts the effective SNR. Meanwhile, the NMSE changes mildly with $P_{\text{B}}$, indicating that with the optimal scaling $\iota^{\star}$, the NMSE is dominated by spatial shape mismatch rather than global amplitude. Across the entire power range, “Rot-BS & Rot-RIS” achieves the best performance in both $\mathcal{U}$ and $R$, together with the smallest NMSE. This gain comes from the additional geometric DoF provided by jointly rotating the active and passive arrays, which reshapes the array manifolds and effectively enlarges the feasible set of beampatterns under the same electronic beamforming resources. Moreover, the RIS-assisted schemes significantly outperform the no-RIS baselines, confirming that the additional passive aperture not only enhances the downlink channels but also provides additional spatial controllability for sensing beampattern shaping.
Directional elements ($b=2$): Figs. 3, 5, and 7 show that the benefit of joint rotation becomes substantially more pronounced when practical directional element patterns are employed. In this case, misalignment between the array boresight and the desired propagation directions can severely reduce the effective channel gain (and thus increase NMSE), potentially driving the weighted utility to negative values for fixed-orientation baselines. By contrast, “Rot-BS & Rot-RIS” maintains the highest utility and sum rate and achieves the lowest NMSE, since it can simultaneously align the BS and RIS boresights with the dominant user/target directions and avoid low-gain regions. These results highlight a key LAWN-specific insight: mechanical geometry control is most valuable when the system operates with directional arrays and/or becomes interference-limited, where purely electronic beamforming alone is insufficient to reliably meet both communication and sensing objectives.
Figure 8: Weighted ISAC utility the number of BS antennas under $b^{\text{R}}=b^{\text{B}}=0$ with $\rho=10$, $N=6\times 6$, $M_{\text{col}}=1$, $P=30$ dBm, and $K=3$.
Figure 9: Sum rate versus the number of BS antennas under $b^{\text{R}}=b^{\text{B}}=0$ with $\rho=10$, $N=6\times 6$, $M_{\text{col}}=1$, $P=30$ dBm, and $K=3$.
Figure 10: Beampattern NMSE versus the number of BS antennas under $b^{\text{R}}=b^{\text{B}}=0$ with $\rho=10$, $N=6\times 6$, $M_{\text{col}}=1$, $P=30$ dBm, and $K=3$.
IV-C Impact of BS Array Size
Figs. 8-10 study the impact of the BS array size by increasing $M_{\text{row}}$ for a $1\times M_{\text{row}}$ BS array under $K=3$ and $P_{\text{B}}=30$ dBm (with $b=0$). As $M_{\text{row}}$ increases, the weighted utility and sum rate (Figs. 8-9) increase steadily because a larger aperture provides higher array gain and more spatial DoF to separate users while shaping the sensing beampattern. In addition, the NMSE decreases with $M_{\text{row}}$ (Fig. 10), showing that more antennas enable a finer beampattern approximation for the same desired template. Notably, the proposed “Rot-BS & Rot-RIS” scheme consistently achieves the best utility-rate-NMSE trade-off across all $M_{\text{row}}$, while the partially rotatable and fully fixed schemes exhibit higher NMSE floors. This indicates that array scaling and geometry adaptation are complementary: increasing $M$ expands the electronic beamforming space, while rotation expands the geometric manifold of feasible array responses, and the combination yields the most robust ISAC performance.
Figure 11: Weighted ISAC utility versus the number of communication users under $b^{\text{R}}=b^{\text{B}}=2$ with $\rho=10$, $N=6\times 6$, $M=2\times 4$, and $P=30$ dBm.
Figure 12: Sum rate versus the number of communication users under $b^{\text{R}}=b^{\text{B}}=2$ with $\rho=10$, $N=6\times 6$, $M=2\times 4$, and $P=30$ dBm.
Figure 13: Beampattern NMSE versus the number of communication users under $b^{\text{R}}=b^{\text{B}}=2$ with $\rho=10$, $N=6\times 6$, $M=2\times 4$, and $P=30$ dBm.
IV-D Impact of User Loading
Figs. 11-13 investigate user scaling under a $2\times 4$ BS array, $N=6\times 6$, $P_{\text{B}}=30$ dBm, and directional elements ($b=2$). As the number of users $K$ increases, the sum rate increases due to multiuser spatial multiplexing gain (Fig. 12), and consequently the weighted utility also increases (Fig. 11). However, Fig. 13 shows that the NMSE generally increases with $K$, which reflects the intrinsic ISAC tension: serving more users requires allocating more spatial resources to interference management and data transmission, leaving less flexibility for beampattern shaping. Importantly, “Rot-BS & Rot-RIS” consistently yields the highest utility and sum rate while achieving the lowest NMSE across all $K$. This demonstrates that joint BS and RIS rotation is particularly beneficial in heavily loaded regimes, where geometric adaptation provides additional leverage to mitigate interference and preserve sensing beampattern fidelity under directional element constraints.
Figure 14: Weighted ISAC utility versus the maximum rotation range $r$ (degrees) under $b^{\text{R}}=b^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 2$, and $P_{\text{B}}=30$ dBm.
Figure 15: Sum rate versus the maximum rotation range $r$ under $b^{\text{R}}=b^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 2$, and $P_{\text{B}}=30$ dBm.
Figure 16: Beampattern NMSE versus the maximum rotation range $r$ (degrees) under $b^{\text{R}}=b^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 2$, and $P_{\text{B}}=30$ dBm.
IV-E Impact of Rotation Range Constraints
Figs. 14-16 evaluate the effect of mechanical rotation limits under $b=2$ with $M=2\times 2$, $N=6\times 6$, and $P_{\text{B}}=30$ dBm. For ease of evaluation, we let each Euler angle, in (24d) and (24e), of both arrays sharing the same rotation range $r$ defined by $r_{\mathrm{i},\min}^{\text{A}}=-r,r_{\mathrm{i},\max}^{\text{A}}=r,\forall\text{A}\in\left\{\text{B},\text{R}\right\},\forall\mathrm{i}\in\{\mathrm{x},\mathrm{y},\mathrm{z}\}$. As the maximum allowable rotation range $r$ increases, the utility and sum rate improve (Figs. 14-15), and the NMSE decreases (Fig. 16), since a larger rotation range provides more geometric DoF to align array boresights and reshape the array response. We further observe diminishing returns: a moderate rotation capability already captures most of the gain, and further increasing $r$ yields only marginal improvements. Moreover, allowing the BS to rotate is generally more influential than rotating the RIS alone, because BS rotation directly impacts both the direct BS-user channels and the BS-RIS illumination, while RIS rotation mainly affects the reflected link. These results suggest that in practical designs, equipping the BS with moderate rotation capability and combining it with even limited RIS rotation can offer a near-optimal rate-sensing trade-off with manageable mechanical complexity.
[…]
IV-F Pareto Frontier of the Rate-Sensing trade-off
Fig. 17 plots the achievable rate-sensing trade-off by sweeping the weight $\rho$ in the weighted-sum objective of (24a). Each marker corresponds to the converged solution of (P1) for a given $\rho$, where we sweep $\rho$ on a logarithmic-like grid $\rho\in\{0.1,0.2,\dots,0.9,1,2,$ $\dots,9,10,20,\dots,90,100,200,\dots,1000\}$. The horizontal axis reports the beampattern NMSE in (22), where a smaller NMSE indicates a more accurate match to the desired sensing beampattern. Hence, moving rightward along each curve corresponds to increasing $\rho$ and placing a stronger emphasis on sensing fidelity, at the cost of reduced downlink sum rate.
Several key observations can be drawn. First, all schemes exhibit a clear and monotonic rate-NMSE trade-off: as $\rho$ increases, the obtained NMSE decreases, but the sum rate gradually degrades. Notably, the curves display a “knee” behavior, where moderate sensing improvement can be achieved with a relatively small rate loss, while pushing NMSE further down incurs a disproportionately large throughput sacrifice. This knee region offers a practical operating point for low-altitude ISAC deployment that balances sensing fidelity and multiuser throughput. Second, the fully rotatable architecture “Rot-BS & Rot-RIS” yields the outermost boundary. Intuitively, BS rotation provides a global geometric DoF that simultaneously impacts the direct BS-user links and the BS-side steering of the cascaded BS-RIS-user paths, thereby improving user separation and interference management. RIS rotation further enriches the controllable spatial response of the reflected link and enables additional beampattern shaping capability, which becomes particularly valuable when stringent sensing fidelity is required. Third, comparing the partially rotatable cases reveals the roles of active versus passive rotations. “Rot-BS & Fix-RIS” achieves a markedly better frontier than “Fix-BS & Rot-RIS”, indicating that BS rotation is more influential in this setup as it directly affects all communication and sensing steering vectors. Nevertheless, enabling RIS rotation still provides a significant frontier expansion beyond “Rot-BS & Fix-RIS”, especially in the sensing-centric regime, where the passive array can help match the desired beampattern with less distortion to the multiuser downlink beams.
@@ V Conclusion
This paper investigated a low-altitude ISAC downlink architecture that exploits mechanical rotations at both the active BS array and the passive RIS to introduce additional spatial DoF beyond conventional electronic beamforming. By incorporating rotation-dependent array geometry and element directivity into a unified directional channel model, we formulated a joint ISAC design that maximizes a weighted utility comprising the downlink sum rate and the sensing beampattern similarity, measured by NMSE. To solve the highly nonconvex problem with coupled variables, we developed an AO framework that iteratively updates the BS transmit beamforming, RIS phase shifts, and the BS and RIS array rotation angles. Specifically, the transmit beamforming subproblem is addressed by combining a QT for the sum-rate term with an MM surrogate for the sensing NMSE, yielding a structured QCQP with an efficient closed-form update under a single power constraint. The RIS phase update is performed on the complex unit-modulus manifold using RCG, while the rotation variables are updated using a projected gradient method with analytically derived Jacobians. Each block update is designed to monotonically improve the objective, and the overall procedure converges to a stationary point of the alternating iterations. Numerical results demonstrated that enabling rotations at both the BS and the RIS consistently enlarges the achievable ISAC performance region compared with fixed-array or partially rotatable baselines, with the gains becoming more pronounced under practical directional element patterns. Moreover, by sweeping the utility weight, we obtained achievable rate-sensing Pareto frontiers that explicitly quantify how mechanical rotations expand the trade-off boundary.
Several important extensions remain for future work. First, incorporating imperfect CSI and sensing-target uncertainty would enable robust designs aligned with practical estimation errors. Second, the time-scale mismatch and actuation overhead of mechanical rotations (e.g., latency, energy consumption, and motion constraints) should be modeled to capture dynamic operation. Finally, extending the framework to wideband channels, multi-target sensing, and discrete or quantized RIS hardware constraints would further strengthen the applicability of rotatable active-passive arrays in realistic low-altitude ISAC systems.